# FusionPortableV2: A Unified Multi-Sensor Dataset for Generalized SLAM Across Diverse Platforms and Scalable Environments

### v1

@@ 5.2 Challenging Factors
Prior to data collection, we acknowledge that practical factors contribute to sensor degradation and potential algorithmic failure. Our data sequences, integrated with the platforms described, aim to comprehensively evaluate algorithm performance in terms of accuracy, efficiency, and robustness. Additionally, we anticipate these sequences will potentially draw the development of novel algorithms.
@@ 5.2.2 Richness of Texture and Structure:
Structured environments can mainly be explained using geometric primitives (e.g., offices or buildings), while semi-structured environments have both geometric and complex elements like trees and sundries. Scenarios like narrow corridors are structured but may cause state estimators. Additionally, texture-rich scenes facilitate visual algorithms to extract stable features (e.g., points and lines), while texture-less may negatively affect the performance. Also, in texture-less environments, only a small amount of events are triggered.
@@ 5.2.3 Dynamic Objects:
Figure 9: Projected point cloud onto the left frame image with our SDK. Points’ color indicate the relative distance. This involves basic implementation including data loader, calibration loader, point cloud manipulation, and camera model.
@@ 5.2.4 Intermittent GNSS:
5.2.5 Scale Variability:
Developing SLAM and perception algorithms for large-scale environments may encounter challenges such as an increased computational load and a heightened risk of perceptual aliasing. The former necessitates stricter demands on algorithm latency and memory usage, whereas the latter requires more accurate long-term associations for place recognition (Lowry et al., 2015), given the potential for environments to include geographically distant yet visually similar locations.
5.3 Sequence Description
@@
Figure 12: GT RGB point cloud map of the (a) campus ($\approx 0.36km^{2}$) and (b) the underground parking lot ($\approx 0.037km^{2}$) with $4mm$-resolution. It almost encompasses the range of most sequences except for those related to vehicles. For detailed map information, please refer to our video presentation.
@@ 5.4 Dataset Organization
Fig. 8 outlines our dataset’s organization. Sensor data were captured using the ROS bag tool. To facilitate download, ROS bags were compressed with 7-Zip. Each bag follows the naming convention $<$platform_env$>$. Sensor calibration parameters are saved in yaml files corresponding to their coordinate frames, such as frame_cam00.yaml for the left camera and frame_cam01.yaml for the right. Given the $6$-month dataset construction period, calibration was performed and documented multiple times, with parameters organized by calibration date (e.g., 20230426_calib). Sequences must utilize the appropriate calibration files and such correspondences are provided in the development package. GT poses at the TUM format are recorded in files matching the sequence names, detailing timestamp, orientation (as Hamilton quaternion), and translation vector per line. GT map at $8cm$ resolution is provided as the pcd format, naming as $<$map_env$>$. Please note that all the data provided have undergone additional post-processing steps, following the procedures detailed in Section 6.
@@ 5.5.1 Software Development Kit (SDK):
We present a Python-only SDK that is both extensible and user-friendly. The kit includes foundational functions such as loading calibration parameters and visualizing them using a TF tree, parsing ROS messages into discrete files, data post-processing, and basic data manipulation). Fig. 9 shows the point cloud projection function provided by the package.
@@ 6.1 Privacy Management
Data collection in public spaces such as the campus and urban roads was conducted with strict adherence to privacy regulations. We employed the anonymization technique introduced in (Burnett et al., 2023) to obscure all human faces and license plates in images from our stereo frame cameras. Building upon the original implementation, we enhanced the algorithm’s efficiency using the ONNX Runtime Library [^8] https://onnxruntime.ai. This upgraded version is now ROS-compatible, offering a valuable resource to the community, and is included in our development tools.
@@ 6.2 GT Data Processing
6.2.1 3-DoF GT Poses of MS60:
The preprocessing initiates with temporal alignment, crucial for synchronizing Leica MS60 total station measurements with sensor data, following the approach proposed in (Nguyen et al., 2022). This synchronization finds the optimal time offset ($\delta$) that minimizes the Absolute Trajectory Error (ATE) between the MS60606060’s recorded poses $\mathcal{T}_{ms60}$ and the SLAM-generated poses $\mathcal{T}_{alg}$. This is done by enumerating offsets at intervals of $0.01s$ to generate various versions of time-shifted poses from the total station poses, as shown in Fig. 13. Upon adjusting their timestamps with the optimal $\delta$, all MS60606060 poses (recorded at $5$-$8Hz$) are resampled to a denser $20Hz$ sequence via cubic spline interpolation. This method not only yields a smooth, continuous trajectory but also synchronizes the data with SLAM algorithm outputs temporally. To maintain interpolated data accuracy, intervals longer than $1s$, potentially signifying obstructions, are omitted from the sequence.
6.2.2 6-DoF GT Poses of 3DM-GQ7:
Each $6$-DoF pose provided by the INS is accompanied by a variance value, indicating the measurement’s uncertainty. This uncertainty increases when the GNSS signal is obstructed. We provide the original data along with tools designed to remove data points with excessive covariance, ultimately converting them into the TUM format. For sequences where intermittent GNSS signals disrupt the INS odometry filter convergence (denoted as sequences vehicle_street00 and vehicle_tunnel00), we recommend using the original GNSS data sampled at $2Hz$, providing 3-DoF GT poses instead.
6.2.3 Quality Analysis of GT Maps:
In constructing our GT maps, we utilized a high-precision Leica scanner for both indoor and outdoor environments. The resulting consolidated RGB point clouds comprise $289$ scanning stations and $342$ inter-scan links, as depicted in Fig. 12. We leveraged Leica Cyclone software for quality assessment of the merged GT maps, resulting in an average error of less than $15mm$ across all pairwise scans, with $90\%$ of these scans having an error of $10mm$ or less. These findings confirm that the precision of GT map far exceeds that of LiDAR-inertial odometry (typically $30mm$), validating its use for evaluation purposes.
@@ 7 Experiment
We manually select $8$ representative sequences ($2$ sequences from each platform) from the dataset to conduct a series algorithm evaluation and verification. Experiments include localization, mapping, and monocular depth estimation.
@@ 7.1.1 Experiment Setting:
As one of the main applications, this dataset can be used to benchmark SOTA SLAM algorithms. Here in, for evaluation of robust localization systems with different input modalities, we carefully select four SOTA SLAM algorithms (including a learning-based method): DROID-SLAM (left frame camera) (Teed and Deng, 2021), VINS-Fusion(LC) (IMU+stereo frame cameras+loop closure) (Qin et al., 2018), FAST-LIO2 (IMU+LiDAR) (Xu et al., 2022), and R3LIVE (IMU+LiDAR+left frame camera) (Lin and Zhang, 2022). The customized data loaders of each evaluated method will be publicly released to foster research.
@@
Table 5: Localization accuracy: we calculate translation ATE $[m]$ for each sequence.
@@
handheld_room00000000 $\bm{0.057}$ $0.058$ $0.063$ $0.118$
handheld_escalator00000000 $0.093$ $\bm{0.085}$ $0.258$ $4.427$
legged_grass00000000 $\bm{0.069}$ $0.327$ $1.801$ $7.011$
legged_room00000000 $\bm{0.068}$ $0.093$ $0.149$ $0.135$
ugv_campus00000000 $\bm{1.486}$ $1.617$ $1.866$ $43.869$
ugv_parking00000000 $0.424$ $\bm{0.271}$ $2.400$ $2.019$
vehicle_campus00000000 $10.070$ $\bm{8.584}$ $66.428$$\times$
@@
Marching from traditional to deep learning-based SLAM methods, we evaluate DROID-SLAM, an end-to-end deep visual SLAM algorithm. We employ DROID-SLAM on the monocular image stream with a pre-trained model [^9] https://github.com/princeton-vl/DROID-SLAM without fine-tuning to present a fair comparison to model-based methods. Meanwhile, testing the limits of its generalization ability is crucial for SLAM. The pre-trained model is trained by supervision from optical flow and poses on the synthetic dataset TartanAir (Wang et al., 2020), covering various conditions (e.g., appearance and viewpoint changes) and environments (e.g., from small-scale indoor to large-scale suburban). All the experiments are conducted on NVIDIA GPU GeForce RTX 3090 with a downsampled image resolution of $320\times 240$. The average runtime is $16$ FPS with a global bundle adjustment (BA) layer. The average GPU memory consumption is below $11$ GB.
@@ 7.1.2 Evaluation:
We choose the typical evaluation metric: mean ATE to evaluate the accuracy of estimated trajectories against the GT using the EVO package. Table 5 reports the quantitative localization results.
In our evaluation of SOTA SLAM systems, we observed that each system’s performance is closely tied to its sensor configuration and algorithmic approach, as outlined in our experimental settings. The analysis, reflected in Table 5, reveals insights into the adaptability and precision of these systems across various environments. Due to the precise geometric information inherent in LiDAR raw data, methods incorporating LiDAR generally exhibit higher accuracy. However, as scene scale increases and becomes more complex (like the highway), segments lacking visual texture or structural features become challenging. FAST-LIO2, which utilizes IMU and LiDAR data, showcased robust performance across a diverse array of environments. This highlights the inherent strength of LiDAR-based systems in tackling various and complex scenarios. In contrast, R3LIVE, which integrates IMU, LiDAR, and visual data, consistently demonstrated superior accuracy in different settings, particularly outperforming FAST-LIO2 in scenarios where LiDAR degradation and jerky motion pattern are present (e.g., ugv_campus00, legged_grass00). However, in environments featuring intricate visual features such as water surfaces or reflective glass on the ugv_parking00, the presence of visual mechanisms in R3LIVE may lead to a performance decrease.
For vision-based methods, VINS-Fusion outperforms DROID-SLAM on average, demonstrating robustness and generalization ability over learning-based methods. In the meantime, it is worth noting that DROID-SLAM with only monocular input beats VINS-Fusion on three sequences: handheld_room00, legged_room00, and ugv_parking00. Compared with the rest of the sequences, these three sequences are of smaller scale with constraint boundaries (i.e., a closed environment). Nevertheless, this result presents promising potential for employing deep learning in SLAM algorithms. With the recent advent of visual foundation models (Oquab et al., 2023), we might be able to unleash the power of data-driven methods to build an accurate and general SLAM foundation model.
These findings emphasize the importance of sensor diversity and well-designed fusion techniques in SLAM systems to address the variability and complexity of real-world environments.
@@
7.2 Evaluation of Mapping
Localization and mapping represent the foundational tasks for robotic navigation, and evaluating trajectory accuracy alone does not suffice to encapsulate the efficacy of such processes comprehensively. Within the framework of SLAM algorithms predicated on Gaussian models, the map serves as an estimand, and its accuracy assessment indirectly mirrors the precision of localization. For the broader spectrum of mapping tasks, whether conducted online or offline, sparse or dense, direct evaluation of map accuracy remains crucial. Hence, a module dedicated to assessing map accuracy has been developed to address this need, ensuring a holistic appraisal of navigational competencies.
7.2.1 Experiment Setting:
For GT map acquisition, high-precision dense RGB point cloud maps were collected using a Leica scanner, with point clouds sampled using a voxel grid $8cm$. Point cloud maps of these sequences were estimated using a SLAM algorithm, with single frame map point clouds sampled at a $10cm$ voxel grid and the aggregated point cloud map of stitched single frames sampled at a $1cm$ grid. Initial transformations for aligning the estimated maps to the GT maps were obtained using CloudCompare. Finally, a correspondence threshold of $20cm$ was established to quantify error metrics, facilitating a rigorous assessment of map accuracy and fidelity.
7.2.2 Evaluation:
@@
Figure 17: This figure shows the unsupervised depth prediction results of our baseline method generalized over different system configurations.
@@
Modals Test Sequences ARD SRD RNSE-linear $[m,\downarrow],$ RNSE-log $\delta<1.25\ [\%,\uparrow]$ $\delta<1.25^{2}\ [\%,\uparrow]$ $\delta<1.25^{3}\ [\%,\uparrow]$
@@
We follow (Hu et al., 2024) to use these metrics for mapping evaluation, as a complement to localization evaluation. In the presence of high-precision RGB point cloud map ground truth, the accuracy of the maps reconstructed by the algorithm can be evaluated. We register the estimated point cloud map $\mathcal{M}$, reconstructed by the algorithm, to the ground-truth point cloud map $\mathcal{G}$, subsequently obtaining the corresponding set of associated points. By computation of precision metrics on this set of associated points, we can also mitigate minor map variations due to temporal discrepancies in ground-truth point cloud collection and slight interference from dynamic obstacles. We compare these two maps based on four metrics: Reconstruction Error (RE) in terms of the Root Mean Squared Error (RMSE), Completeness (COM), and Chamfer Distance (CD). They are defined as below:
  • Reconstruction Error computes the average point-to-point distance between $\mathcal{M}$ and $\mathcal{G}$ (Pan et al., 2022):
    $\textit{RE}=\sqrt{\frac{1}{|\mathcal{M}|}\sum_{\bm{p}\in\mathcal{M}}{\underbrace{\min(\tau,\ \|\bm{p}-\bm{q}\|)}_{d(\bm{p},\mathcal{G})}}^{2}},$ (1)
  • Completeness is equal to $\mathrm{COM}=|\mathcal{G}^{\prime}|/|\mathcal{G}|$ that describes how does $\mathcal{M}$ cover extent of $\mathcal{G}$. $\mathcal{G}^{\prime}$ is the subset of $\mathcal{G}$. Each element of $\mathcal{G}^{\prime}$ has a nearby point from $\mathcal{M}$ such as
    $\textit{COM}=\{\mathbf{q}\in\mathcal{G}^{\prime}\ |\ \ \ \exists\mathbf{p}\in\mathcal{M},||\mathbf{p}-\mathbf{q}||\leq\tau\}.$ (2)
Fig. 15 employs the map evaluation module to present the assessment outcomes of the FAST-LIO algorithm across six indoor and outdoor data sequences, with varying colors representing the accuracy levels across different map regions. The map accuracy estimated by FAST-LIO notably decreases in outdoor large-scale scenes (Fig. 15 (a)) or areas with dense vegetation (Fig. 15 (c) and Fig. 15 (f)), attributable to significant measurement noise from trees or overall z-axis drift in outdoor LiDAR Odometry and Mapping applications. Conversely, indoor settings, barring the effects introduced by dynamic obstacles (Fig. 15 (f)), predominantly exhibit high map quality (Fig. 15 (e)).
Table 6 further delineates the map evaluation results for F2L and R3L across six indoor and outdoor sequences, with R3L retaining only points with RGB colors, hence showing inferior performance on the COM metric compared to F2L. However, F2L outperforms R3L in most scenes in terms of RE and CD metrics, particularly in handheld and quadruped robot sequences. In expansive campus environments, both algorithms exhibit comparable CD metrics, as seen in scenarios like handheld_escalator and ugv_campus. Fig. 16 illustrates the mapping results of R3LIVE on the handheld_room00 sequence, employing a color scheme consistent with that of Fig. 15. The presence of glass within the room introduces noise, resulting in some blurred regions within the map. This depiction underscores the impact of environmental features on mapping clarity.

### v2

@@ 5.2 Challenging Factors
Prior to data collection, we acknowledge that practical factors contribute to sensor degradation and potential algorithmic failure. Our data sequences, integrated with the platforms described, aim to comprehensively evaluate algorithm performance in terms of accuracy, efficiency, and robustness. Additionally, we anticipate these sequences will draw the development of novel algorithms.
@@ 5.2.2 Richness of Texture and Structure:
Structured environments (e.g., offices or buildings) can mainly be explained using geometric primitives, while semi-structured environments have both geometric and complex elements like trees and sundries. Scenarios like narrow corridors are structured but may challenge state estimators. Additionally, texture-rich scenes facilitate visual algorithms to extract stable features (e.g., points and lines), while texture-less may negatively affect the performance. Also, in texture-less environments, only a small amount of events is triggered.
@@ 5.2.3 Dynamic Objects:
Figure 9: The projected point cloud onto the left frame image using our SDK shows points’ colors indicating relative distances. This involves a basic implementation, including a data loader, calibration loader, point cloud manipulation, and camera model.
@@ 5.2.4 Intermittent GNSS:
5.2.5 Scale Variability:
Developing SLAM and perception algorithms for large-scale environments may encounter challenges such as an increased computational load and a heightened risk of perceptual aliasing. The former necessitates stricter demands on algorithm latency and memory usage, whereas the latter requires more accurate long-term associations for place recognition (Yin et al., 2024), given the potential for environments to include geographically distant yet visually similar locations.
5.2.6 Viewpoint Change:
We consider the viewpoint change from two perspectives: “yaw-change” and “roll-pitch-change”. The former often occurs in sequences with loops. Several sequences in our dataset feature at least one loop, posing challenges for place recognition and image matching algorithms. For instance: ugv_transition00 and ugv_transition01 contain loops in both indoor and outdoor environments; vehicle_multilayer00 captures multiple loops across different floors of a multi-story parking lot; and vehicle_campus00 and vehicle_campus01 collectively cover the entire HKUST campus, where loops are also present. We also provide a similar sequence (campus_road_day) in FusionPortable (Jiao et al., 2022), collected $15$ months ago, during which several buildings and roads were updated.
The latter aspect is particularly relevant to the cross-view localization (CVL) problem (Shi et al., 2023), which involves viewpoint changes between down-facing satellite images and ground-level images. Open-source satellite images can be obtained from tools such as Google Earth. Since several sequences include accurate GT geo-localization information and covers diverse scenarios (e.g., campus, urban roads, downhill mountain roads), it can serve as a challenging benchmark for CVL.
5.3 Sequence Description
@@
Figure 12: GT RGB point cloud map of the (a) campus ($\approx 0.36km^{2}$) with $8cm$-resolution and (b) the underground parking lot ($\approx 0.037km^{2}$) with $4mm$-resolution. It almost encompasses the range of most sequences except for those related to vehicles. For detailed map information, please refer to our video presentation.
@@ 5.4 Dataset Organization
Fig. 8 outlines our dataset’s organization. Sensor data were captured using the ROS bag tool1212endnote: 12https://wiki.ros.org/rosbag. ROS bags are used due to their numerous advantages, such as mature tools for debugging, visualization of the tf tree, especially with multiple sensors present, a broadcasting mechanism for batching output messages of different types, and the ability to convert them into individual files. To facilitate download, ROS bags were compressed with 7-Zip. Each bag follows the naming convention $<$platform_environment$>$. Sensor calibration parameters are saved in yaml files naming as $<$frame_id$>$ (e.g., frame_cam00.yaml for the left RGB camera and frame_cam01.yaml for the right.) During the $6$-month dataset construction period, calibration was performed and documented multiple times, with parameters organized by calibration date (e.g., 20230426_calib). Sequences must utilize the appropriate calibration files and such correspondences are provided in the development package. GT poses at the TUM format are recorded in files matching the sequence names, detailing timestamp, orientation (as Hamilton quaternion), and translation vector per line. GT map is provided as the pcd format, naming as $<$location_environment$>$. Please note that all the data provided have undergone additional post-processing steps, following the procedures detailed in Section 6.
@@ 5.5.1 Software Development Kit (SDK):
We present a Python-only SDK that is both extensible and user-friendly. The kit includes foundational functions such as loading calibration parameters and visualizing them using a TF tree, parsing ROS messages into discrete files, data post-processing, and basic data manipulation. Fig. 9 shows the point cloud projection function provided by the package.
@@ 6.1 Privacy Management
Data collection in public spaces such as the campus and urban roads was conducted with strict adherence to privacy regulations. We employed the anonymization technique1313endnote: 13https://github.com/understand-ai/anonymizer introduced in (Burnett et al., 2023) to obscure all human faces and license plates in images from our stereo frame cameras. Building upon the original implementation, we enhanced the algorithm’s efficiency using the ONNX Runtime Library1414endnote: 14https://onnxruntime.ai. This upgraded version is now ROS-compatible, offering a valuable resource to the community, and is included in our development tools.
@@ 6.2 GT Data Processing
6.2.1 3-DoF GT Poses of Total Station:
The preprocessing initiates with temporal alignment, crucial for synchronizing Leica MS60606060 total station measurements with sensor data, following the approach proposed in (Nguyen et al., 2022). This synchronization finds the optimal time offset that minimizes the Absolute Trajectory Error (ATE) between the MS60606060’s recorded poses $\mathcal{T}_{ms60}$ and the SLAM-generated poses $\mathcal{T}_{alg}$. This is done by enumerating offsets at intervals of $0.01s$ to generate various versions of time-shifted poses from the total station poses, as shown in Fig. 13. Upon adjusting their timestamps with the optimal $\delta$, all MS60606060 poses (recorded at $5$-$8Hz$) are resampled to a denser $20Hz$ sequence via cubic spline interpolation. This method not only yields a smooth, continuous trajectory but also temporally synchronizes the data with SLAM algorithm outputs. To maintain data accuracy, intervals longer than $1s$ indicating obstructions are omitted from the sequence.
6.2.2 6-DoF GT Poses of INS:
Each $6$-DoF pose provided by the INS is accompanied by a variance value, indicating the measurement’s uncertainty. This uncertainty increases when the GNSS signal is obstructed. For a fair comparison, we manually removed data points with excessive uncertainty to serve as ground truth. We also provide the original data along with relevant tools for users to process the data according to their needs. For the sequences vehicle_street00 and vehicle_tunnel00, where intermittent GNSS signals disrupt the INS odometry filter convergence, we directly used the original GNSS data sampled at $2Hz$, providing 3-DoF GT poses instead. For the sequence vehicle_multilayer00, collected in a multi-layer parking garage, most of the trajectory is indoors without access to GNSS signals. Therefore, we used the results from the FAST-LIO2 algorithm as the baseline, given the rich structured information available. The start and end points of the sequence are located on the rooftop, where GNSS data can be used as loop closure references.
6.2.3 Accuacy of GT Maps:
In the construction of our ground truth (GT) maps, we employed a high-precision Leica scanner to capture detailed environments both indoors and outdoors. Fig. 12 displays the complete RGB point cloud map of the entire campus and a section of the underground parking garage. The fine details observable in the depicted areas highlight the high quality of both the point cloud and its color fidelity. According to the Leica Cyclone software report, the GT map exhibits an average error of less than $15mm$ across all pairwise scans, the average error is $3.4mm$ and with $90\%$ of these scans maintaining an error margin of $10mm$ or less. The precision of our GT map exceeds that of point cloud maps constructed via current LiDAR SLAM technologies by nearly two orders of magnitude, making it suitable for algorithm evaluation.
@@ 7 Experiment
We select $8$ representative sequences ($2$ sequences from each platform) from the dataset to conduct a series algorithm evaluation and verification. Experiments include localization, mapping, and monocular depth estimation.
@@ 7.1.1 Experiment Setting:
As one of the main applications, this dataset can be used to benchmark SOTA SLAM algorithms. Here in, for evaluation of localization systems with different input modalities, we select four SOTA SLAM algorithms (including a learning-based method): DROID-SLAM (left frame camera) (Teed and Deng, 2021), VINS-Fusion (LC) (IMU + stereo frame cameras, with loop closure enabled) (Qin et al., 2018), FAST-LIO2 (IMU+LiDAR) (Xu et al., 2022), and R3LIVE (IMU+LiDAR+left frame camera) (Lin and Zhang, 2022). The customized data loaders of each method are publicly released to foster research.
@@
Marching from traditional to deep learning-based SLAM methods, we evaluate DROID-SLAM, an end-to-end deep visual SLAM algorithm. We employ DROID-SLAM on the monocular image stream with a pre-trained model1515endnote: 15https://github.com/princeton-vl/DROID-SLAM without fine-tuning to present a fair comparison to model-based methods. Meanwhile, testing the limits of its generalization ability is crucial for SLAM. The pre-trained model is trained by supervision from optical flow and poses on the synthetic dataset TartanAir (Wang et al., 2020), covering various conditions (e.g., appearance and viewpoint changes) and environments (e.g., from small-scale indoor to large-scale suburban). All the experiments are conducted on NVIDIA GPU GeForce RTX $3090$ with a downsampled image resolution of $320\times 240$. The average runtime is $16$ FPS with a global bundle adjustment (BA) layer. The average GPU memory consumption is below $11$ GB.
7.1.2 Evaluation:
We choose the typical evaluation metric: mean ATE to evaluate the accuracy of estimated trajectories against the GT using the EVO package.1616endnote: 16https://github.com/MichaelGrupp/evo Table 5 reports the quantitative localization results.
Our evaluation of SOTA SLAM systems, as summarized in Table 5, demonstrates that each system’s performance varies across different environments, depending on its sensor configuration and algorithmic approach. Due to the precise geometric information inherent in LiDAR raw data, methods incorporating LiDAR generally exhibit higher accuracy. However, as scene scale increases and becomes more complex (like the highway), segments lacking visual texture or structural features become challenging. FAST-LIO2, which utilizes IMU and LiDAR data, showcased robust performance across a diverse array of environments. This highlights the inherent strength of LiDAR-based systems in tackling various and complex scenarios. In contrast, R3LIVE, which integrates IMU, LiDAR, and visual data, consistently demonstrated superior accuracy in different settings, particularly outperforming FAST-LIO2 in scenarios where LiDAR degradation and jerky motion pattern are present (e.g., ugv_campus00, legged_grass00). However, in environments featuring intricate visual features such as water surfaces or reflective glass on the ugv_parking00, the presence of visual mechanisms in R3LIVE may lead to a performance decrease.
For vision-based methods, VINS-Fusion outperforms DROID-SLAM on average, demonstrating robustness and generalization ability over learning-based methods. However, it is important to note that DROID-SLAM, using only monocular input, surpasses VINS-Fusion in three specific sequences: legged_room00, and ugv_parking00. These sequences are characterized by smaller-scale environments with constrained boundaries, such as closed rooms or parking areas. This result highlights the promising potential of employing deep learning techniques in SLAM algorithms, particularly in scenarios where the environment is more confined and structured. The superior performance of DROID-SLAM in these specific cases suggests that learning-based methods can excel in certain conditions, despite their overall lower average performance compared to traditional approaches like VINS-Fusion.
Table 5: Localization accuracy: we calculate translation ATE $[m]$ for each sequence. The best result among all methods is shown in bold, and the best result among vision-based methods (VINS-Fusion and DROID-SLAM) is underlined.
@@ 7.1.2 Evaluation:
handheld_room00000000 $\bm{0.057}$ $0.058$ $\underline{0.063}$ $0.118$
handheld_escalator00000000 $0.093$ $\bm{0.085}$ $\underline{0.258}$ $4.427$
legged_grass00000000 $\bm{0.069}$ $0.327$ $\underline{1.801}$ $7.011$
legged_room00000000 $\bm{0.068}$ $0.093$ $0.149$ $\underline{0.135}$
ugv_campus00000000 $\bm{1.486}$ $1.617$ $\underline{1.866}$ $43.869$
ugv_parking00000000 $0.424$ $\bm{0.271}$ $2.400$ $\underline{2.019}$
vehicle_campus00000000 $10.070$ $\bm{8.584}$ $\underline{66.428}$$\times$
@@
7.2 Evaluation of Mapping
Localization and mapping represent the foundational tasks for robotic navigation, and evaluating trajectory accuracy alone does not suffice to encapsulate the efficacy of such processes comprehensively. Within the framework of SLAM algorithms predicated on Gaussian models, the map serves as a crucial output, and its accuracy assessment indirectly mirrors the precision of localization. For the broader spectrum of mapping tasks, whether conducted online or offline, sparse or dense, direct evaluation of map accuracy remains crucial. Hence, a module dedicated to assessing map accuracy has been developed to address this need, ensuring a holistic appraisal of navigational competencies.
7.2.1 Experiment Setting:
For evaluating point cloud maps estimated by SOTA SLAM algorithms, we first downsampled the estimated maps using a 0.1m grid. Initial alignment with the GT map was performed using CloudCompare software. We set the maximum threshold distance for corresponding points at $0.2m$. Thus, after evaluation, point pairs with a distance less than $0.2m$ were considered as the same point for distance calculation.
@@ 7.2.2 Evaluation:
We use the mapping evaluation metrics in PALoc (Hu et al., 2024b) to complement our localization evaluation. After initial alignment, the error metrics were then calculated using our map evaluation library as introduced in Section 3.3.1. In the presence of high-precision RGB point cloud map ground truth, the accuracy of the maps reconstructed by the algorithm can be evaluated. We register the estimated point cloud map $\mathcal{M}$, reconstructed by the algorithm, to the ground-truth point cloud map $\mathcal{G}$, subsequently obtaining the corresponding set of associated points. By computation of precision metrics on this set of associated points, we can also mitigate minor map variations due to temporal discrepancies in ground-truth point cloud collection and slight interference from dynamic obstacles. We compare these two maps based on four metrics: Reconstruction Error (RE) in terms of the Root Mean Squared Error (RMSE), Completeness (COM), and Chamfer Distance (CD). They are defined as below:
  • Reconstruction Error computes the average point-to-point distance between $\mathcal{M}$ and $\mathcal{G}$ (Pan et al., 2024):
    $\textit{RE}=\sqrt{\frac{1}{|\mathcal{M}|}\sum_{\bm{p}\in\mathcal{M}}{\underbrace{\min(\tau,\ \|\bm{p}-\bm{q}\|)}_{d(\bm{p},\mathcal{G})}}^{2}}.$ (1)
  • Completeness describes how does $\mathcal{M}$ cover extent of $\mathcal{G}$. $\mathcal{G}^{\prime}$ is the subset of $\mathcal{G}$. Each element of $\mathcal{G}^{\prime}$ has a nearby point from $\mathcal{M}$ such as
    $\begin{split}\textit{COM}&=\frac{|\mathcal{G}^{\prime}|}{|\mathcal{G}|},\\ \mathcal{G}^{\prime}&=\{\mathbf{q}\in\mathcal{G}\ |\ \ \ \exists\mathbf{p}\in\mathcal{M},||\mathbf{p}-\mathbf{q}||\leq\tau\}.\end{split}$ (2)
Fig. 15 employs the map evaluation module to present the assessment outcomes of the FAST-LIO2 algorithm across six indoor and outdoor data sequences, with varying colors representing the accuracy levels across different map regions. The map accuracy estimated by FAST-LIO2 notably decreases in outdoor large-scale scenes (Fig. 15 (a)) or areas with dense vegetation (Fig. 15 (c) and Fig. 15 (f)), attributable to significant measurement noise from trees or overall z-axis drift in outdoor LiDAR odometry and mapping applications. Conversely, indoor settings, barring the effects introduced by dynamic obstacles (Fig. 15 (f)), predominantly exhibit high map quality (Fig. 15 (e)).
Table 6 further delineates the map evaluation results for FAST-LIO2 and R3LIVE across $6$ indoor and outdoor sequences, with R3LIVE retaining only points with RGB colors, hence showing inferior performance on the COM metric compared to FAST-LIO2. However, FAST-LIO2 outperforms R3LIVE in most scenes in terms of RE and CD metrics, particularly in handheld and quadruped robot sequences. In expansive campus environments, both algorithms exhibit comparable CD metrics, as seen in scenarios like handheld_escalator and ugv_campus. Fig. 16 illustrates the mapping results of R3LIVE on the handheld_room00 sequence, employing a color scheme consistent with that of Fig. 15. The presence of glass within the room introduces noise, resulting in some blurred regions within the map. This depiction underscores the impact of environmental features on mapping clarity.
@@
Figure 17: Results of the unsupervised depth prediction method Liu et al. (2023) which can generalize over different environments.
@@
Model Test Sequence ARD SRD RMSE-linear $[m,\downarrow],$ RMSE-log $\delta<1.25\ [\%,\uparrow]$ $\delta<1.25^{2}\ [\%,\uparrow]$ $\delta<1.25^{3}\ [\%,\uparrow]$